OpenStudy (anonymous):

x+1/5=2x-2/6 help me solve this equation?!

OpenStudy (unklerhaukus):

is it x+(1/5) = 2x-(2/6) or (x+1)/5 = (2x-2)/6 ?

OpenStudy (anonymous):

(x+1)/5 = (2x-2)/6

OpenStudy (shamim):

2x-x=1/5+2/6 right?

OpenStudy (unklerhaukus):

(x+1)/5 = (2x-2)/6 start by multiplying both sides of the equation by 5 and then multiply both sides by 6 what do you get ?

OpenStudy (anonymous):

Just the equation because what I by multiplying 5 and 6 on both sides didn't make sense

OpenStudy (unklerhaukus):

(x+1)/5 = (2x-2)/6 multiplying both sides of the equation by 5 5 * (x+1)/5 = 5 * (2x-2)/6 (x+1) = 5 * (2x-2)/6

OpenStudy (unklerhaukus):

can you now multiply both sides by 6 @Ilovedogs21 ?

OpenStudy (anonymous):

I got 4 for the final answer

OpenStudy (unklerhaukus):

if you want to check you answer x=4 sub it into the original equation (x+1)/5 = (2x-2)/6 (4+1)/5 = (2(4)-2)/6 If this equation is true, then you have found the correct solution